The Bigbee Bottling Company is contemplating the replacement of one of its bottling machines with a newer and more efficient one. The old machine has a book value of $625,000 and a remaining useful life of 5 years. The firm does not expect to realize any return from scrapping the old machine in 5 years, but it can sell it now to another firm in the industry for $265,000. The old machine is being depreciated by $125,000 per year, using the straight-line method.

The new machine has a purchase price of $1,125,000, an estimated useful life and MACRS class life of 5 years, and an estimated salvage value of $140,000. The applicable depreciation rates are 20%, 32%, 19%, 12%, 11%, and 6%. It is expected to economize on electric power usage, labor, and repair costs, as well as to reduce the number of defective bottles. In total, an annual savings of $225,000 will be realized if the new machine is installed. The company's marginal tax rate is 35%, and it has a 12% WACC.

  1. What initial cash outlay is required for the new machine? Round your answer to the nearest dollar. Negative amount should be indicated by a minus sign.
    $
  2. Calculate the annual depreciation allowances for both machines and compute the change in the annual depreciation expense if the replacement is made. Round your answers to the nearest dollar.
    Year Depreciation Allowance, New Depreciation Allowance, Old Change in Depreciation
    1 $ $ $
    2
    3
    4
    5
  3. What are the incremental net cash flows in Years 1 through 5? Round your answers to the nearest dollar.
    Year 1 Year 2 Year 3 Year 4 Year 5
    $ $ $ $ $
  4. Should the firm purchase the new machine?
    -Select-Yes or No

    Support your answer. The input in the box below will not be graded, but may be reviewed and considered by your instructor.
  5. In general, how would each of the following factors affect the investment decision, and how should each be treated?
    1. The expected life of the existing machine decreases.

    The input in the box below will not be graded, but may be reviewed and considered by your instructor.

    2. The WACC is not constant, but is increasing as Bigbee adds more projects into its capital budget for the year.

    The input in the box below will not be graded, but may be reviewed and considered by your instructor.
Answer #1

Part b)

Year Dep Allow New Dep Allow Old Change in depreciation (Dep Allow New-Old)
1 (1,125,000-140,000)*20% = 197,000 125,000 72,000
2 (1,125,000-140,000)*32% = 315,200 125,000 190,200
3 (1,125,000-140,000)*19% = 187,150 125,000 62,150
4 (1,125,000-140,000)*12% = 118,200 125,000 -6,800
5 (1,125,000-140,000)*11% = 108,350 125,000 -16,650

Part a, c & d)

Sl.No Year 0 1 2 3 4 5
Initial Outlay
i Purchase of new machine (Given) -1,125,000
ii Sale of old machine (Given) 265,000
iii Net initial outlay (i+ii) (part A) -860,000
Incremental net cash flows
iv Annual savings if new machine is installed (Given) 225,000 225,000 225,000 225,000 225,000
v Change in depreciation (refer part b) 72,000 190,200 62,150 -6,800 -16,650
vi Incremental profit before tax (iv-v) 153,000 34,800 162,850 231,800 241,650
vii Taxes @ 35% (vi*35%) 53,550 12,180 56,998 81,130 84,578
viii Incremental profit after tax (vi-vii) 99,450 22,620 105,853 150,670 157,073
ix Add back: change in depreciation (v) 72,000 190,200 62,150 -6,800 -16,650
x Incremental net cash flows (viii+ix) (Part c) 171,450 212,820 168,003 143,870 140,423
Terminal cash flows
xi Salvage value of new machine (Given) 140,000
xii Net cash flows (iii+x+xi) -860,000 171,450 212,820 168,003 143,870 280,423
xiii PVF @ 12% 1/[1.12^year] 1 0.8929 0.7972 0.7118 0.6355 0.5674
xiv Discounted cash flows (xii*xiii) -860,000 153,088 169,660 119,584 91,429 159,112

Part d) Net present value of replacement = ΣDiscounted cash flows = -167,127

Since NPV is negative, so the firm should not purchase new machine.